Telescope and AGw Commissioning at LFBG with IRTC - 25-26 April 2009

Observers: A. Rakich, M. Pedani, J. Hill, J. Heidt (Tucson)
Telescope Operator: A. Ceranski
Software Support: (phone) Telescope Support:

Summary

Entire night was lost to high winds.

Preparation

Details

Appendix

HBS Oil has been running for about an hour, and is in relatively steady state since the dome is closed.

ECS PLC Readings

18:56 MST ECS-HBS Temperature Setpoint = -0.24 degC, ECS-Coolant(Oil) Return Temperature +0.46 degC


Air------------Steel--------------
C-Ring Upper 8.52 7.76
C-Ring Middle 8.19 7.65
C-Ring Lower 8.23 7.76
Extension Left 8.54 8.12
Extension Right 7.48 7.88
WB Upper 7.64 7.67
WB Middle Left 7.68 7.70
WB Middle Right 7.74 7.76
WB Lower Right 8.17 8.43
Cell Lower Right 7.78 7.61

HBS PLC Readings from MCSPU

Oil Reservoir 2.8
Input Cooling 5.5
Output Cooling 1.5
Ta 6.1

Elevation Tin 8.7 - 9.3
Elevation Tout 8.1 - 9.4
Azimuth Tin 7.8 - 9.2
Azimuth Tout 8.4 - 10.3

Fluke Camera Readings (uncorrected)

LR Dino Egg 5.1
Oily Side of LR EL axial pad 6.4
Exterior of Left C-Ring 4.3
WB Center Joint 4.3
C-Ring Right Lower (near Tim's sensor) 4.1
RR Lateral Pad 6.0
DX Mirror Center hole (glass w/o aluminum) 3.3
RF HBS Input Hoses 3.6
Oil in HBS Tank -1.4

Note that the Fluke Thermal Imager seems to be off by -2 degC. While the Ranger ST non-contact meter seems well calibrated when looking at an ice bath. Later I checked the Fluke Thermal Imager against the ice bath and it seemed OK.

PMC Readings

DX Glass Therms 6.96
DX Ambient Therms 7.40
SX Glass Therms 6.57
SX Ambient Therms 7.68

20:31 Changed HBS Thermal Loop Offset from -8.0 degC to -9.5 degC in lbt.conf.

21:10 Stop and restart ECS for the change to take effect.

-- JohnHill - 26 Apr 2009
Topic revision: r2 - 26 Apr 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback